# Capillary $L_{p}$-Christoffel-Minkowski problem

### v1

@@ 1. Introduction
A natural question is how this picture changes in the presence of a boundary. In the capillary setting, one considers hypersurfaces in the half-space that meet a fixed supporting hyperplane at a prescribed contact angle $\theta\in(0,\pi/2)$. For the top-order case $k=n$, capillary versions of the $L_{p}$-Minkowski problem have been developed in a series of recent works. For $p\geq 1$, Mei, Wang and Weng solved the capillary $L_{p}$-Minkowski problem via the continuity method in [MWW25a, MWW25b]. For (n+1)<p<1-(n+1)<p<1, even solutions were constructed in [HI25] by means of an iterative scheme based on the curvature image operator, and a unified curvature flow approach was later introduced in [HHI25], treating the even capillary $L_{p}$-Minkowski problem for all p>(n+1)p>-(n+1).
For $k<n$, the capillary analogue of the Christoffel–Minkowski problem prescribes $\sigma_{k}(\tau^{\sharp}[s])$ on the capillary spherical cap $\mathcal{C}_{\theta}$ and couples the interior equation with a Robin boundary condition encoding the contact angle. This capillary analogue was solved in [HIS25], where the existence of smooth, strictly convex, $\theta$-capillary hypersurfaces was established under conditions on the prescribed function that are tailored to the applicability of a constant rank theorem.
@@ 2. Preliminaries
(2.21) $V(s_{0},\dots,s_{k},\underbrace{\ell,\ldots,\ell}_{n-k\ \text{times}})=\frac{1}{n+1}\int_{\mathcal{C}_{\theta}}s_{0}Q_{k}\bigl(\tau^{\sharp}[s_{1}],\dots,\tau^{\sharp}[s_{k}],\underbrace{\tau^{\sharp}[\ell],\dots,\tau^{\sharp}[\ell]}_{n-k\ \text{times}}\bigr).$
@@
(2.22) $V(s_{0},\underbrace{s,\dots,s}_{k\ \text{times}},\ell,\dots,\ell)=\frac{1}{n+1}\int_{\mathcal{C}_{\theta}}s_{0}\frac{\sigma_{k}\bigl(\tau^{\sharp}[s]\bigr)}{\binom{n}{k}}$.
Figure 1. Capillary vs. classical outer parallel hypersurfaces
@@
Therefore $\nu(x^{\prime})=\nu(x)$. Since $\Sigma$ is strictly convex, the Gauss map $\nu:\Sigma\to\mathbb{S}_{\theta}^{n}$ is injective, hence $x^{\prime}=x$.
@@
Since $\Sigma_{t}$ has the same capillary support function as $\Sigma+t\mathcal{C}_{\theta}$, we conclude
@@
For $\rho>0$ and a Borel set $\omega\subset\mathcal{C}_{\theta}$ we define the local outer parallel set of $\widehat{\Sigma}$ in the directions of $\omega$ by
@@
Let
@@
Let $\theta\in(0,\pi/2)$ and suppose $\Sigma\subset\mathbb{R}^{n+1}_{+}$ is a strictly convex, $\theta$-capillary hypersurface. For a Borel set $\omega\subset\mathcal{C}_{\theta}$ we define the capillary $k$-th area measure of $\widehat{\Sigma}$ over $\omega$ by
@@ 3.1. Rotationally symmetric hypersurfaces
where we used that $C_{0}>1$ and $q\in[1,p]$. Since $\Sigma$ is rotationally symmetric, $r_{\mathrm{in}}=r_{\mathrm{out}}$ and thus $s\leq r_{\mathrm{in}}+H$. Now, by 3.3 and 2.8,
@@ 6. Existence and uniqueness
$=(n+1)\binom{n}{k}V\bigl(s_{2},\underbrace{s_{1},\ldots,s_{1}}_{k\text{ copies}},\ell,\ldots,\ell\bigr)$
$\geq(n+1)\binom{n}{k}V\bigl(\underbrace{s_{1},\ldots,s_{1}}_{k+1\text{ copies}},\ell,\ldots,\ell\bigr)^{\frac{k}{k+1}}V\bigl(\underbrace{s_{2},\ldots,s_{2}}_{k+1\text{ copies}},\ell,\ldots,\ell\bigr)^{\frac{1}{k+1}}$
@@
is uniformly elliptic with $C^{\alpha}$ coefficients. Define the boundary operator

### v2

@@ 1. Introduction
A natural question is how this picture changes in the presence of a boundary. In the capillary setting, one considers hypersurfaces in the half-space that meet a fixed supporting hyperplane at a prescribed contact angle $\theta\in(0,\pi/2)$. For the top-order case $k=n$, capillary versions of the $L_{p}$-Minkowski problem have been developed in a series of recent works. For $p\geq 1$, Mei, Wang and Weng solved the capillary $L_{p}$-Minkowski problem via the continuity method in [MWW25a, MWW25c]. For (n+1)<p<1-(n+1)<p<1, even solutions were constructed in [HI25] by means of an iterative scheme based on the curvature image operator, and a unified curvature flow approach was later introduced in [HHI25], treating the even capillary $L_{p}$-Minkowski problem for all p>(n+1)p>-(n+1).
For $k<n$, the capillary analogue of the Christoffel–Minkowski problem prescribes $\sigma_{k}(\tau^{\sharp}[s])$ on the capillary spherical cap $\mathcal{C}_{\theta}$ and couples the interior equation with a Robin boundary condition encoding the contact angle. This capillary analogue was solved in [HIS25], where the existence of smooth, strictly convex, $\theta$-capillary hypersurfaces was established under conditions on the prescribed function that are tailored to the applicability of a constant rank theorem. The existence of a solution to the capillary Christoffel-Minkowski problem was also established in [MWW25c], subject to an additional assumption concerning the existence of a suitable homotopy path.
@@ 2. Preliminaries
(2.21) $V(s_{0},\dots,s_{k},\underbrace{\ell,\ldots,\ell}_{(n-k)-\text{times}})=\frac{1}{n+1}\int_{\mathcal{C}_{\theta}}s_{0}Q_{k}\bigl(\tau^{\sharp}[s_{1}],\dots,\tau^{\sharp}[s_{k}],\underbrace{\tau^{\sharp}[\ell],\dots,\tau^{\sharp}[\ell]}_{(n-k)-\text{times}}\bigr).$
@@
(2.22) $V(s_{0},\underbrace{s,\dots,s}_{k-\text{times}},\ell,\dots,\ell)=\frac{1}{n+1}\int_{\mathcal{C}_{\theta}}s_{0}\frac{\sigma_{k}\bigl(\tau^{\sharp}[s]\bigr)}{\binom{n}{k}}$.
@@
Thus $\nu(x^{\prime})=\nu(x)$. Since $\Sigma$ is strictly convex, the Gauss map $\nu:\Sigma\to\mathbb{S}_{\theta}^{n}$ is injective, hence $x^{\prime}=x$.
@@
Since $\Sigma_{t}$ has the same capillary support function as $\Sigma+t\mathcal{C}_{\theta}$, we conclude that
@@
Remark 2.5.
The notion of capillary outer parallel sets for the capillary convex bodies was first introduced in [MWW25c], while the relation $\Sigma_{t}=\Sigma+t\mathcal{C}_{\theta}$ was observed in [MWWX25, Rem. 2.17]. 2.4 clarifies the connection between capillary and classical outer parallel hypersurfaces, see Figure 1.
Figure 1. Capillary vs. classical outer parallel hypersurfaces
For $\rho>0$ and a Borel set $\omega\subset\mathcal{C}_{\theta}$, the local outer parallel set of $\widehat{\Sigma}$ in the directions of $\omega$ can be defined by
@@
The local Steiner-type formula was previously stated in [MWW25c]. For completeness, we give a proof here. Let
@@
Let $\theta\in(0,\pi/2)$ and suppose $\Sigma\subset\mathbb{R}^{n+1}_{+}$ is a strictly convex, $\theta$-capillary hypersurface. For a Borel set $\omega\subset\mathcal{C}_{\theta}$, the capillary $k$-th area measure of $\widehat{\Sigma}$ over $\omega$ can be defined by (see also [MWW25c])
@@ 3.1. Rotationally symmetric hypersurfaces
where we used that $C_{0}>1$ and $q\in[1,p]$. Since $\Sigma$ is rotationally symmetric, $r_{\mathrm{in}}=r_{\mathrm{out}}$ and thus $s\leq r_{\mathrm{in}}+H$. Now, by 3.3 and 2.9,
@@ 6. Existence and uniqueness
$=(n+1)\binom{n}{k}V\bigl(s_{2},\underbrace{s_{1},\ldots,s_{1}}_{k-\text{times}},\ell,\ldots,\ell\bigr)$
$\geq(n+1)\binom{n}{k}V\bigl(\underbrace{s_{1},\ldots,s_{1}}_{(k+1)-\text{times}},\ell,\ldots,\ell\bigr)^{\frac{k}{k+1}}V\bigl(\underbrace{s_{2},\ldots,s_{2}}_{(k+1)-\text{times}},\ell,\ldots,\ell\bigr)^{\frac{1}{k+1}}$
@@
is uniformly elliptic with $C^{\infty}$ coefficients. Define the boundary operator